# Are automated proof assistants ready for semigroup research? Orientation-preserving mappings and proof assistant Lean

### v1

@@ Abstract
An orientation-preserving mapping is not always defined by how it acts on triples of elements. Although this fact is simple and well-known, it sometimes gets overlooked, resulting in wrong statements in publications. Automated proof assistants are a technology that is supposed to ensure that proofs exactly match the results, thus pre-empting mistakes. In this paper we successfully formalise the definition of orientation-preserving mappings in proof assistant software Lean and construct a computer-verified proof of the above fact.
@@ 1 Introduction
Order-preserving mappings are defined by how they act on pairs of elements of the chain. Orientation-preserving mappings are ‘almost’ defined by how they act on triples of elements of the chain, but not quite; there are exceptions. Relevant publications where this distinction is studied are Levi and Mitchell (2006) [levimitchell2006], Higgins and Vernitski (2022) [higgins2022orientation], Fernandes and Vernitski [fernandes2025groups]. Although it is well known that orientation-preserving mappings are ‘not quite’ defined by how they act on triples of elements of the chain, this fact frequently gets overlooked and has led to small mistakes in some published results; indeed, paper Higgins and Vernitski (2022) [higgins2022orientation] required a corrigendum Higgins and Vernitski (2024) [higgins2024correction], and paper Fernandes (2023) [fernandes2023CKMS] required a corrigendum Fernandes (2024) [fernandes2024CKMS].
Automated proof assistants are a kind of technology that ensures that proofs exactly match the results, thus pre-empting mistakes in proofs. Are automated proof assistants sufficiently developed to verify facts like the ones described above? In this paper we successfully formalise the definition of orientation-preserving mappings in proof assistant software Lean and construct a computer-verified proof of the fact that orientation-preserving mappings are not always defined by how they act on triples of elements of the chain.
@@ 2 Orientation-preserving mappings
Fix a positive integer $n$. The set $0,1,\dots,n-1$} is denoted by $[n]$. By a mapping we mean a transformation on $[n]$. For simpler notation, a mapping $0\mapsto a_{0},1\mapsto a_{1},\dots,n-1\mapsto a_{n-1}$ will be denoted by $(a_{0},a_{1},\dots,a_{n-1})$.
The standard definition of an orientation-preserving mapping is as follows, see, for example, [higgins2022orientation, fernandes2023CKMS, fernandes2025groups]. Let $s=(a_{1},a_{2},\dots,a_{t})$ be a sequence of $t$ elements of $[n]$. We say that $s$ is cyclic [respectively, anti-cyclic] if there is no more than one index $i\in\{1,\dots,t\}$ such that $a_{i}>a_{i+1}$ [respectively, $a_{i}<a_{i+1}$], where $a_{t+1}$ denotes $a_{1}$.
Given a mapping $f$ whose domain is $\{a_{1}<\dots<a_{t}\}$, we say that $f$ is orientation-preserving [respectively, orientation-reversing] if the sequence of its images $(f(a_{1}),f(a_{2}),\dots,f(a_{t}))$ is cyclic [respectively, anti-cyclic].
@@
Draw $[n]$ around a circle clockwise. Inside that circle, draw another copy of $[n]$ around a circle, also clockwise. Then draw connectors induced by the mapping. If the connectors can be drawn in such a way that they do not intersect, the mapping is orientation-preserving; for example, mapping $(1,2,3,0)$ shown in Figure 1 on the left, is orientation-preserving. If the connectors intersect in whichever way you lay them out, the mapping is not orientation-preserving; for example, mapping $(0,1,0,1)$ shown in Figure 1 on the right, is not orientation-preserving.
To describe the concept of an orientation-reversing mapping, draw $[n]$ around a circle clockwise, and inside that circle, draw another copy of $[n]$ around a circle, but anti-clockwise. Then draw connectors induced by the mapping. If the connectors can be drawn in such a way that they do not intersect, the mapping is orientation-reversing. If the connectors intersect in whichever way you lay them out, the mapping is not orientation-reversing.
Generally speaking, we have observed that making conclusions concerning planar diagrams is tricky and frequently leads to mistakes. Famous examples include the controversy regarding the first proof of the Jordan’s curve theorem and Kempe’s incorrect proof of the four-colour theorem. Examples known to us from our own recent practice include [lisitsa2023describing] and [howie2020planar] correcting errors in earlier papers of other authors. In semigroup theory, orientation-preserving mappings present an example of a confusing type of planar diagrams, as evidenced by the need for corrigenda [higgins2024correction, fernandes2024CKMS]. This is why this concept and, to be specific, the mapping $(0,1,0,1)$ seems to us a suitable test case for applying an automated proof assistant.
@@ 3.1 What are Lean and mathlib?
Lean 4 [lean4] is the current major version of the Lean programming language. In the context of mathematical research, Lean is often referred to as a proof assistant or an interactive theorem prover (though these terms may give an inflated impression of its current capabilities in the production of proofs). The primary current use of Lean in this context is in proof verification: the researcher will attempt to formalise a theorem in Lean, i.e. provide a fully elaborated proof with completely explicit, checkable steps which can be formally verified by Lean’s typechecking logic. If the verification fails, Lean will attempt to provide useful error messages for the researcher.
mathlib4 is the current version of mathlib, a project maintained by the Lean community [mathlib], and is a library of formalised mathematics built in Lean. It contains a very large base of definitions and theorems, across most areas of undergraduate mathematics, which the researcher can build on and adapt. As a well-written repository, it is also an excellent (if somewhat advanced) resource for learning how to formalise theorems in Lean; the more beginner-friendly interactive online tutorial book Mathematics in Lean [mil] is built on top of mathlib.
@@ 3.3 Inductive types
3.4 The code: calculating that $(0,1,0,1)$ is unoriented
The following represents our attempt to formalise this small result in Lean 4: we do not claim that this is the most efficient or idiomatic way to do so. (In fact, we will see in §3.5 below that it is not well-suited to all purposes.) All code below was developed and tested in Lean 4.25.0 with mathlib 4.25.0.
@@ 3.4 The code: calculating that $(0,1,0,1)$ is unoriented
The Lean parser responds with orientationSort.none, confirming that this sequence is neither cyclic nor anticyclic.
3.5 Next steps: proving semigroup theorems
So far, most of our work has been definitions; a sensible next step from here would be to formalise a small theorem, such as [higgins2022orientation, Theorem 3] (after the corrigendum of [higgins2024correction]), which says that the orientation of a sequence of rank $\neq 2$ is determined by the orientations of its subsequences of length 3. It is usually a good idea to break a theorem down into the smallest possible lemmas: for instance, the user may begin by attempting to prove that a subsequence of a cyclic sequence is cyclic, or even that a subsequence of an increasing sequence is increasing. Here, we use the built-in definition Sublist.
@@ 3.6 Adopting Lean in research and teaching
    For those who are not experienced in software development, installing Lean and its toolchain, keeping it updated to the latest version, and so on, can already be a nontrivial barrier. For researchers looking to try out Lean without installing it, we recommend the Lean playground [playground], a browser-based implementation of Lean with no installation requirements.
@@
    There are many excellent community-written books aimed at Lean beginners with no knowledge of type theory, of which we mention only [mil] and [tpil]. The Lean Game Server [gameserver], a repository of interactive games aimed at teaching Lean syntax, can also serve as a good first introduction.
@@
    Lean 4 and mathlib4 are still in active development, and at time of writing, new versions are released once a month. Most changes add new functionality, and the occasional deprecation of functionality is minor, especially for new users: for instance, InChain (mentioned above) is new as of September 2025, and the older proposition Chain is now deprecated. This is unlikely to break old code (within reasonable timeframes), but keeping up with the latest versions of Lean 4 and mathlib4 will ensure that the user has access to the largest, most up-to-date bank of definitions and lemmas.
@@
Another possible application of Lean we can see is in teaching, especially among students who have some basic familiarity with programming: an introductory course in logic or proofs may benefit from the high standards of formalisation and the immediate feedback from Lean’s proof checker. Using mathlib, this could also draw on the students’ previous knowledge of abstract algebra (proving that, for instance, $(ab)^{-1}=b^{-1}a^{-1}$ in a group is an easy but nontrivial task in Lean), elementary number theory or discrete mathematics. The book [mil] is a good source of inspiration here.

### v2

# Capturing properties of planar diagrams in Lean proof assistant software

@@ Abstract
Automated proof assistants are a technology pre-empting mistakes in mathematics. In our practice we have seen that reasoning about planar diagrams is difficult to both humans and computers. One example that has led to wrong statements in publications is that an orientation-preserving mapping is not always defined by how it acts on triples of elements. In this paper we formalise orientation-preserving mappings in proof assistant software Lean and report on our take-aways.
Keywords. Automated proof, proof assistant, theorem prover.
@@ 1 Introduction
In our practice we have observed that making conclusions concerning planar diagrams is tricky and frequently leads to mistakes. Famous examples include the controversy regarding the first proof of the Jordan curve theorem and Kempe’s incorrect proof of the four-colour theorem. Examples known to us from our own recent practice include [LLV23] and [HW20] correcting errors in earlier papers of other authors.
The example considered in this article is also taken from our recent practice. Order-preserving mappings (which is how monotonic functions are called in semigroup theory) are defined by how they act on pairs of elements of the chain, that is, a mapping $f$ is called order-preserving if from $a\leq b$ it follows that $f(a)\leq f(b)$. Orientation-preserving mappings are a generalization of order-preserving mappings. As we will see in the following section, orientation-preserving mappings are ‘almost’ defined by how they act on triples of elements of the chain, but not quite; there are exceptions. Relevant publications where this distinction is studied are [LM06, HV22, FV25]. Although it is well known that orientation-preserving mappings are ‘not quite’ defined by how they act on triples of elements of the chain, this fact frequently gets overlooked and has led to small mistakes in some published results; indeed, paper [HV22] required a corrigendum [HV24], and paper [FER23] required a corrigendum [FER24].
Automated proof assistants are a kind of technology that ensures that proofs exactly match the intended results, thus pre-empting mistakes. At the same time, it is difficult to write code for an automated proof assistant that can be used to reason about planar diagrams. Are automated proof assistants sufficiently developed to verify facts like the ones described above? In this paper we formalise orientation-preserving mappings in proof assistant software Lean and report on our take-aways.
Note that within this article, it is not our purpose to define planarity as such. For our purposes, planarity is a sweet spot where both humans and computers, for different reasons, find it difficult to make correct inferences, and therefore, this is where we can draw examples showing how human reasoning and automated proof assistants can complement one another.
@@ 2 Orientation-preserving mappings
Fix a positive integer $n$. The set $\{0,1,\dots,n-1\}$ is denoted by $[n]$. By a mapping we mean a transformation on $[n]$. For simpler notation, a mapping $0\mapsto a_{0},1\mapsto a_{1},\dots,n-1\mapsto a_{n-1}$ will be denoted by $(a_{0},a_{1},\dots,a_{n-1})$.
The standard definition of an orientation-preserving mapping is as follows, see, for example, [HV22, FER23, FV25]. Let $s=(a_{1},a_{2},\dots,a_{t})$ be a sequence of $t$ elements of $[n]$. We say that $s$ is cyclic [respectively, anti-cyclic] if there is no more than one index $i\in\{1,\dots,t\}$ such that $a_{i}>a_{i+1}$ [respectively, $a_{i}<a_{i+1}$], where $a_{t+1}$ denotes $a_{1}$.
Given a mapping $f$ whose domain is $\{a_{1}<\dots<a_{t}\}$, we say that $f$ is orientation-preserving (respectively, orientation-reversing) if the sequence of its images $(f(a_{1}),f(a_{2}),\dots,f(a_{t}))$ is cyclic (respectively, anti-cyclic).
@@
Draw $[n]$ around a circle clockwise. Around that circle, draw another copy of $[n]$ around a circle, also clockwise. Then draw connectors induced by the mapping. If the connectors can be drawn in such a way that they do not intersect, the mapping is orientation-preserving; for example, mapping $(1,2,3,0)$ shown in Figure 1 on the left, is orientation-preserving. If the connectors intersect in whichever way you lay them out, the mapping is not orientation-preserving; for example, mapping $(0,1,0,1)$ shown in Figure 1 on the right, is not orientation-preserving. Mapping $(0,1,0,1)$ is an example demonstrating that an orientation-preserving mapping is not always defined by how it acts on triples of elements; indeed, every restriction of mapping $(0,1,0,1)$ to $3$ elements is orientation-preserving. As we discussed in the previous section, this counterexample, although well-known, was overlooked in the statements of some results in [HV22, FER23].
To describe the concept of an orientation-reversing mapping, draw $[n]$ around a circle clockwise, and around that circle, draw another copy of $[n]$ around a circle, but anti-clockwise. Then draw connectors induced by the mapping. If the connectors can be drawn in such a way that they do not intersect, the mapping is orientation-reversing. If the connectors intersect in whichever way you lay them out, the mapping is not orientation-reversing.
In the following sections we use proof assistant software Lean to demonstrate that sequence $(0,1,0,1)$ is neither cyclic nor anti-cyclic (we do not expand this result to proving that mapping $(0,1,0,1)$ is neither orientation-preserving nor orientation-reversing because this would distract us from our main topic of making inferences about planar diagrams). We then proceed to considering a theorem [HV22, Theorem 3] stating that if a sequence does not include exactly $2$ distinct elements then its being cyclic (or anti-cyclic) is determined by the orientations of its subsequences of length $3$.
@@ 3.1 What are Lean and mathlib?
Lean 4 [10] is the current major version of the Lean programming language. In the context of mathematical research, Lean is often referred to as a proof assistant or an interactive theorem prover (though these terms may give an inflated impression of its current capabilities in the production of proofs). The primary current use of Lean in this context is in proof verification: the researcher will attempt to formalise a theorem in Lean, i.e. provide a fully elaborated proof with completely explicit, checkable steps which can be formally verified by Lean’s typechecking logic. If the verification fails, Lean will attempt to provide useful error messages for the researcher.
mathlib4 is the current version of mathlib, a project maintained by the Lean community [11], and is a library of formalised mathematics built in Lean. It contains a very large base of definitions and theorems, across most areas of undergraduate mathematics, which the researcher can build on and adapt. As a well-written repository, it is also an excellent (if somewhat advanced) resource for learning how to formalise theorems in Lean; the more beginner-friendly interactive online tutorial book Mathematics in Lean [2] is built on top of mathlib.
@@ 3.3 Inductive types
3.4 The code: $(0,1,0,1)$ is neither cyclic nor anti-cyclic
The following represents our attempt to formalise this small result in Lean 4: we do not claim that this is the most efficient or idiomatic way to do so. (In fact, we will see in §3.5 below that it is not well-suited to all purposes.) All code below was developed and tested in Lean 4.27.0 with mathlib 4.27.0.
@@ 3.4 The code: $(0,1,0,1)$ is neither cyclic nor anti-cyclic
The Lean parser responds with orientationSort.none, confirming that this sequence is neither cyclic nor anti-cyclic.
3.5 Next steps: proving a theorem
So far, most of our work has been definitions; a sensible next step from here would be to formalise a small theorem, such as [HV22, Theorem 3] (after the corrigendum of [HV24]), which says that the orientation of a sequence of rank $\neq 2$ is determined by the orientations of its subsequences of length 3. It is usually a good idea to break a theorem down into the smallest possible lemmas: for instance, the user may begin by attempting to prove that a subsequence of a cyclic sequence is cyclic, or even that a subsequence of an increasing sequence is increasing. Here, we use the built-in definition Sublist.
@@ 3.5 Next steps: proving a theorem
To summarise, on the one hand, in this section we gave an overview of a selection of Lean tools that can be successfully used to make progress towards proving [HV22, Theorem 3], and on the other hand, we saw that actually proving [HV22, Theorem 3] in Lean would require inordinate amount of work, so we did not proceed to writing complete code for this proof.
@@ 3.6 Adopting Lean in research and teaching
    For those who are not experienced in software development, installing Lean and its toolchain, keeping it updated to the latest version, and so on, can already be a nontrivial barrier. For researchers looking to try out Lean without installing it, we recommend the Lean playground [9], a browser-based implementation of Lean with no installation requirements.
@@
    There are many excellent community-written books aimed at Lean beginners with no knowledge of type theory, of which we mention only [2] and [14]. The Lean Game Server [1], a repository of interactive games aimed at teaching Lean syntax, can also serve as a good first introduction.
@@
    Lean 4 and mathlib4 are still in active development, and at time of writing, new versions are released once a month. Most changes add new functionality, and the occasional deprecation of functionality is minor, especially for new users: for instance, IsChain (used above) is new as of September 2025, and the older proposition Chain is now deprecated. This is unlikely to break old code (within reasonable timeframes), but keeping up with the latest versions of Lean 4 and mathlib4 will ensure that the user has access to the largest, most up-to-date bank of definitions and lemmas.
@@
Another possible application of Lean we can see is in teaching, especially among students who have some basic familiarity with programming: an introductory course in logic or proofs may benefit from the high standards of formalisation and the immediate feedback from Lean’s proof checker. Using mathlib, this could also draw on the students’ previous knowledge of abstract algebra (proving that, for instance, $(ab)^{-1}=b^{-1}a^{-1}$ in a group is an easy but nontrivial task in Lean), elementary number theory or discrete mathematics. The book [2] is a good source of inspiration here.